OpenStudy (anonymous):

simplify:

OpenStudy (anonymous):

\[\sqrt[3]{16}\] write using a simplified radical

myininaya (myininaya):

so 16=4*4=2*2*2*2, so the cube root of 16 can be simplified to 2*{cube root of (2)}

OpenStudy (anonymous):

so it woudl look like this?\[\sqrt[2]{2}\]

myininaya (myininaya):

2*(cube root of 2)

myininaya (myininaya):

\[2\sqrt[3]{2}\]